OpenStudy (anonymous):

What is the value of 7P3?

OpenStudy (anonymous):

what is the value of a calculator http://www.wolframalpha.com/input/?i=7P3

OpenStudy (solomonzelman):

\(\Large\color{green}{ 7~P~3 =\frac{7!}{(7-3)!}=\frac{7!}{4!}=\frac{4!\times 5\times 6 \times 7}{4!} =?}\) can you finish ?